a) Price of Pepsi is Rs.120 and its quantity demanded is 300 bottles. Suppose due to Raman Pepsi company reduces price of it to Rs. 108 ,due to it its demand increases to 360 bottles. Calculate elasticity and also mention the type of elasticity.
1
Expert's answer
2020-05-03T16:57:57-0400
"p_1=120; p_2=108"

"Q_1=300; Q_2=360"


"E= \frac{Q_2-Q_1}{p_2-p_1}\times \frac{p_1+p_2}{Q_1+Q_2}"



"E=\frac {360-300}{108-120}\times \frac{120+108}{360+300}=-1.72"


"|E|>1"

Based on the results obtained, the demand for this product is elastic, since the absolute value of the coefficient obtained is more than 1.